OpenStudy (destro):

Divide. Express your answer in lowest terms. 1 4 = 8

OpenStudy (destro):

igreen

OpenStudy (destro):

ash2326

OpenStudy (destro):

thomaster

OpenStudy (anonymous):

Hey, if you want to tag people, you have to have a @ symbol in front of their usernames. :)

OpenStudy (anonymous):

Like this: @ash2326

OpenStudy (anonymous):

And can you write your question more clearly?

OpenStudy (destro):

never mind

OpenStudy (destro):

@igreen

OpenStudy (igreen):

Sorry, Is this it? \(\dfrac{1}{8} \div \dfrac{4}{1}\)

OpenStudy (destro):

no \[\frac{ 1 }{ 8 }\div 4\]

OpenStudy (igreen):

Oh, okay, well dividing is the same thing as multiplying by the reciprocal. So: \(\dfrac{1}{8} \div 4 \rightarrow \dfrac{1}{8} \times \dfrac{1}{4}\) Can you multiply that? @destro